OpenStudy (usukidoll):

Let x be a number on (0, pi/2). Solve the inequality sin x > cos x

OpenStudy (usukidoll):

I haven't dealt with a problem like this in a while... x.x

OpenStudy (anonymous):

look at the graphs, or at the unit circle you can see that they are the same at \(x=\frac{\pi}{4}\)

OpenStudy (usukidoll):

you mean the sin x graph and the cos x graph?

OpenStudy (anonymous):

if \(x<\frac{\pi}{4}\) then cosine is larger, whereas if \(x>\frac{\pi}{4}\) you will see that sine is larger

OpenStudy (usukidoll):